# A Neural Transformer Framework for Simultaneous Tasks of Segmentation, Classification, and Caller Identification of Marmoset Vocalization

### v1

@@
A Neural Transformer Framework for Simultaneous Tasks of Segmentation, Classification, and Caller Identification of Marmoset Vocalization
Bin Wu1    Sakriani Sakti2,1    Shinnosuke Takamichi3,1    Satoshi Nakamura2,1,4
@@
4The Chinese University of Hong Kong, China
@@ 1 Abstract
Marmoset, a highly vocalized primate, has become a popular animal model for studying social-communicative behavior and its underlying mechanism. In the study of vocal communication, it is vital to know the caller identities, call contents, and vocal exchanges. Previous work of a CNN has achieved a joint model for call segmentation, classification, and caller identification for marmoset vocalizations. However, the CNN has limitations in modeling long-range acoustic patterns; the Transformer architecture that has been shown to outperform CNNs, utilizes the self-attention mechanism that efficiently segregates information parallelly over long distances and captures the global structure of marmoset vocalization. We propose using the Transformer to jointly segment and classify the marmoset calls and identify the callers for each vocalization.
@@ 2 Introduction
Researchers have been using the marmoset as an animal model to study diseases and mechanisms related to vocal communication. Uesaka et al. [14] developed a marmoset model of autism disorder — marked by the deficits in social communication and verbal interaction and by verbal preservation — by feeding pregnant mothers with valproic acid. By using the autism model, they aimed to study the developmental vocal characteristics of autism for early diagnosis and investigate potential medicines to relieve the autism symptoms. The ability of vocal communication is shaped by innate and empirical factors. Researchers manipulate autism-related genes of marmosets [6] to study innate gene-function relationships. Researchers manipulate environments including visual or auditory sensory inputs [8] and parental interaction [11] to study the empirical modification of communicative behaviors.
@@
Oikarinen et al. used a dataset of recordings from paired of pairs marmosets housed together in single cages, enabling close-range vocal interaction [7, 9]. The dataset comprises 36 sessions totally 38 hours, with 8 call types: trill, phee, trillphee, twitter, chirp, tsik, ek, and chatter, along with a noise type that indicates the silences between the calls. Applying a CNN model on the dataset, they achieved simultaneous segmentation, classification, and caller identification [9].
However, while previous works as [18] and [9] used an RNN or a CNN that maps acoustic segments to call labels, the Transformer structure has been proven to outperform RNN [15] and CNN [4] in both sequential language processing and high-dimensional vision tasks. Transformer utilizes self-attention mechanism that efficiently segregates information parallelly over long distances and captures the global structure of marmoset vocalization more efficiently than RNN and more effectively than CNN. While Transformer is typically constrained by the quadratic complexity of input token length [15], making it challenging to process high-dimensional input spectra that discriminate the marmoset calls such as a phee and a trill. We can address the limitation using Vision Transformer [4] that patchizes the high-dimensional input. We propose to use the Transformer framework for simultaneous segmentation, classification, and caller identification of marmoset vocalizations.
@@
Figure 1: The two-stream Transformer model. Two Transformer models process two spectral segments extracted from a dual-channel audio recording to map to a vocal call label. The call label of ’phee2’ denotes that the phee call comes from the second marmoset of the pair in the cage; the ’noise’ denotes the non-call segment between calls, used for segmenting calls to non-calls.
@@ 3 Model
We use a two-stream Transformer model (Figure 1) on the dual-audio recordings with two simultaneously recorded channels that come from two interacting animals. The model employs two stream transformer encoders that process sliding spectral segments from each channel to classify into a target label. The target labels indicate call types and caller identities (e.g., the ’tr2’ label denotes a trill call from the second of two animals in interactions). The labels also indicate segmentation information: ’noise’ denotes the non-call segment between calls. This approach allows our two-stream Transformer model to simultaneously perform three crucial tasks:
@@ 5.1 Dataset
Numer of pairs 8 pairs 1 pair 1 pair
@@ 5.3.1 Our CNN backbone model
5.3.2 Our Transformer backbone model
@@
5.3.3 Our CNN/Transformer system with backbone model
@@
To use the system to monitor marmoset behavior for long-hour recordings, we applied a streaming technique to enhance model prediction speed. This process involves two main steps: First, we split the test audios into non-overlapping 2500ms segments. Second, we created 50 sub-segments from each 2500ms segment, using a window size of 500ms and a window shift of 50ms: the first 41 sub-segments are complete within the current segment, while the last 9 concatenate parts from the next 2500ms segment. During prediction, after feeding a batch into the model, we obtain predicted 50ms intervals corresponding to 50 subsegments in the batch. This approach, inspired by [9], allows for efficient processing of long-duration audio recordings.
@@
Table 2: Results of our CNN and Transformer systems and [9]’s best system on the same dataset [7].
F-score Recall Prec. Noise Acc. Call Acc. Total Acc.
Best system of [9] 0.7686 0.7212 0.8227 0.9963 0.7212 0.9899
Our CNN system 0.7901 0.7575 0.8257 0.9962 0.7575 0.9907
Our Transformer system 0.794 0.7576 0.8341 0.9964 0.7576 0.9909
@@ 6 Result and discussion
Our systems can be applied to raw, long-hour audio recordings to simultaneously segment and classify calls, as well as identify callers. This capability provides a valuable tool for recording marmoset interactions, which can support future studies on social behavior, development, and abnormalities in marmoset vocalizations. Such research could offer insights into communication, evolution, and dysfunction of the vocal langauge. Additionally, the Transformer update of our system offers potential for developing a unified multimodal model that integrates both video and audio inputs.
@@ 7 Conclusion
We implemented a two-stream Transformer system for dual-channel audio recordings, enabling simultaneous classification, segmentation, and caller identification. This Transformer system outperforms previous CNN approaches. Our system can efficiently process long-hour or full-day marmoset recordings, providing rich information about vocal communication between marmosets in spontaneous interactions. This capability enhances the use of marmosets as an animal model for studying language evolution, development, and dysfunction in a communication context.

### v3

# A Transformer Model for Segmentation, Classification, and Caller Identification of Marmoset Vocalization

@@
A Transformer Model for Segmentation, Classification, and Caller Identification of Marmoset Vocalization
Bin Wu,1    Shinnosuke Takamichi,1,3    Sakriani Sakti,1,2    Satoshi Nakamura1,2,4
@@
4The Chinese University of Hong Kong, Shenzhen, China
@@ 1 Abstract
Marmoset, a highly vocalized primate, has become a popular animal model for studying social-communicative behavior and its underlying mechanism comparing with human infant linguistic developments. In the study of vocal communication, it is vital to know the caller identities, call contents, and vocal exchanges. Previous work of a CNN has achieved a joint model for call segmentation, classification, and caller identification for marmoset vocalizations. However, the CNN has limitations in modeling long-range acoustic patterns; the Transformer architecture that has been shown to outperform CNNs, utilizes the self-attention mechanism that efficiently segregates information parallelly over long distances and captures the global structure of marmoset vocalization. We propose using the Transformer to jointly segment and classify the marmoset calls and identify the callers for each vocalization.
@@ 2 Introduction
Researchers have been using the marmoset as an animal model to study diseases and mechanisms related to vocal communication comparing with human infant linguistic developments. Uesaka et al. [14] developed a marmoset model of autism disorder — marked by the deficits in social communication, impaired verbal interaction, and verbal perseveration — by feeding pregnant mothers with valproic acid. Using the autism model, they aimed to study the developmental vocal characteristics of autism for early diagnosis and investigate potential medicines to relieve the autism symptoms. The ability of vocal communication is shaped by innate and empirical factors. Researchers manipulate autism-related genes of marmosets [6] to study innate gene-function relationships. Researchers manipulate environments including visual or auditory sensory inputs [8] and parental interaction [11] to study the empirical modification of communicative behaviors.
@@
Oikarinen et al. used a dataset of recordings from paired of pairs marmosets housed together in single cages, enabling close-range vocal interaction [7, 9]. The dataset comprises 36 sessions totally 38 hours, with 8 call types: trill, phee, trillphee, twitter, chirp, tsik, ek, and chatter, along with a noise type that indicates the silences between the calls. Applying a CNN model on the dataset, they achieved segmentation, classification, and caller identification [9].
However, while previous works as [18] and [9] used an RNN or a CNN that maps acoustic segments to call labels, the Transformer structure has been proven to outperform RNN [15] and CNN [4] in both sequential language processing and high-dimensional vision tasks. Transformer utilizes self-attention mechanism that efficiently segregates information parallelly over long distances and captures the global structure of marmoset vocalization more efficiently than RNN and more effectively than CNN. While Transformer is typically constrained by the quadratic complexity of input token length [15], making it challenging to process high-dimensional input spectra that discriminate the marmoset calls such as a phee and a trill. We can address the limitation using Vision Transformer [4] that patchizes the high-dimensional input. We propose to use the Transformer model for segmentation, classification, and caller identification of marmoset vocalizations.
@@
Figure 1: The two-stream Transformer model. Two Transformer models process two spectral segments extracted from a dual-channel audio recording to map to a vocal call label. The call label of ’phee2’ denotes that the phee call comes from the second marmoset of the pair in the cage; the ’noise’ denotes the non-call segment between calls, used for segmenting calls to non-calls. We used two output heads to enable simultaneous call prediction, allowing detection of concurrent calls of the animal pair such as ’trill1’ and ’twitter2’.
@@ 3 Model
We use a two-stream Transformer model (Figure 1) on the dual-audio recordings with two simultaneously recorded channels that come from two interacting animals. The model employs two stream transformer encoders that process sliding spectral segments from each channel to classify into a target label. The target labels indicate call types and caller identities (e.g., the ’tr2’ label denotes a trill call from the second of two animals in interactions). The labels also indicate segmentation information: ’noise’ denotes the non-call segment between calls. This approach allows our two-stream Transformer model to perform three crucial tasks:
@@ 5.1 Dataset
Number of pairs 8 pairs 1 pair 1 pair
@@ 5.3.1 Our CNN backbone model
5.3.2 Proposed Transformer backbone model
@@
5.3.3 Proposed CNN/Transformer system with backbone model
@@
To use the system to monitor marmoset behavior for long-hour recordings, we applied a streaming technique to enhance model prediction speed. This process involves two main steps: First, we split the test audios into non-overlapping 2500ms segments. Second, we created 50 sub-segments from each 2500ms segment, using a window size of 500ms and a window shift of 50ms: the first 41 sub-segments are complete within the current segment, while the last 9 concatenate parts from the next 2500ms segment. During prediction, after feeding a batch into the model, we obtain predicted 50ms intervals corresponding to 50 subsegments in the batch. This approach, inspired by [9], allows for efficient processing of long-duration audio recordings. We also implemented spectral feature extraction with Pytorch that uses GPU for better efficiency.
@@
Table 2: Results of our CNN and proposed Transformer systems and [9]’s best system on the same dataset [7]
Systems F-score Recall Prec.
Best system of [Oikarinen, 2019] 0.7686 0.7212 0.8227
Our CNN system 0.7901 0.7575 0.8257
Proposed Transformer 0.794 0.7576 0.8341
Systems Total Acc. Noise Acc. Call Acc.
Best system of [Oikarinen, 2019] 0.9899 0.9963 0.7212
Our CNN system 0.9907 0.9962 0.7575
Proposed Transformer 0.9909 0.9964 0.7576
@@ 6 Result and discussion
Our systems can be applied to raw, long-hour audio recordings to segment and classify calls, as well as identify callers. This capability provides a valuable tool for recording marmoset interactions, which can support future studies on social behavior, development, and abnormalities in marmoset vocalizations. Such research could offer insights into communication, evolution, and dysfunction of the vocal language of marmoset. The research also helps facilitate comparisons between marmoset and human infant vocal development in family environments. Additionally, the Transformer update of our system offers potential for developing a unified multimodal model that integrates both video and audio inputs.
@@ 7 Conclusion
We implemented a two-stream Transformer system that enables classification, segmentation, and caller identification for marmoset vocalization. This Transformer system outperforms previous CNN approaches. Our system can efficiently process long-hour or full-day marmoset recordings, providing rich information about vocal communication between marmosets in spontaneous interactions. This capability enhances the use of marmosets as an animal model for studying language evolution, development, and dysfunction in a communication context such as the comparative research between vocal development in marmosets and human infants in family environments.